OpenStudy (anonymous):

Factor Completely 1/256w^2+50/64w^3+625/16 please help

OpenStudy (saifoo.khan):

\[\frac{1}{256w^2+50} \div \frac{64w^3+625}{16}\]Do u have soething like that?

OpenStudy (anonymous):

no it is all in a straight line i actually wrote is wrong its.. 1/256w^6-50/64w^3+625/16

OpenStudy (saifoo.khan):

ummm, ok,

OpenStudy (saifoo.khan):

Alright, can u please use brackets?

OpenStudy (anonymous):

{1/256w^6}-{50/64w^3}+{625/16} is this better?

OpenStudy (saifoo.khan):

(1/256w^6)-(50/64w^3)+(625/16) that?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[\Large \frac{10000w^6-200w^3+1}{256w^6}\]

OpenStudy (anonymous):

thank you